# TorchCP: A Python Library for Conformal Prediction

### v4

@@ Appendix E The Advantage of our low-coupling design
Low-coupling design is one of the most fundamental concepts in object-oriented design, which minimizes the dependency a class has on other classes. The low-coupling design of TorchCP ensures flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, most current packages focus on a high-coupling design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or custom score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our low-coupling design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement scores for predictors. Moreover, some libraries (e.g., TorchUQ) implemented various non-conformity scores as functions, but this style suffers from extending to those scores with extra hyperparameters (e.g., the $k$ and $\lambda$ of RAPS).
@@ Appendix F Comparison with related libraries
Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning (DL) scenarios differs significantly, particularly in supporting modern DL models like GNNs and LLMs, as well as their scalability for large-scale datasets. In Table 8, we present a concise comparison of open-source CP libraries, focusing on supported model types and functionalities:
@@
In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with DL applications.
nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression. It lacks GPU acceleration, batch processing, and compatibility with DL models like DNN Classifier, DNN Regressor, GNNs, or LLMs, making it unsuitable for complex DL scenarios. While nonconformist implements predictors and scores as different classes following a low-coupling design, its algorithmic coverage is limited, implementing only inductive CP and aggregated CP with 5 scores.
@@
Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports conformal classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks. However, it lacks GPU acceleration, batch processing, and seamless extensions to modern models like GNNs or LLMs, limiting scalability for engineers tackling large-scale or complex deep learning problems. Similar to TorchUQ, Crepes implements non-conformity scores as functions passed to predictors, so it also suffers from supporting scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS).
PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern DL applications. It does not support advanced DL models such as DNN, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)). PUNCC defines nonconformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.

### v5

@@ Appendix E The Advantage of our low-coupling design
Low-coupling design is one of the most fundamental concepts in object-oriented design, which minimizes the dependency a class has on other classes. The low-coupling design of TorchCP ensures flexibility in integrating diverse training algorithms, non-conformity score functions, and prediction workflows, accommodating the wide variety of CP methods and tasks. In contrast, most current packages focus on a high-coupling design where predictors and scoring functions are often predefined within a specific predictor (e.g., Split CP). This limits extensibility when new CP predictors or user-defined score functions are needed. TorchCP ’s decoupled design allows users to independently select or define new trainers, scores, or predictors, enabling flexible integration of different components. In what follows, we elaborate on the concrete benefits of our low-coupling design:
  • Flexible combination: users can freely combine Trainer, Score, and Predictor modules to customize conformal prediction workflows. For instance, a practitioner can use different non-conformity scores (e.g., LAC or APS) within any predictor or trainer, while other packages only implement specified scores for predictors. Moreover, some libraries (e.g., TorchUQ) implemented various non-conformity scores as functions, but this style suffers from extending to those scores with extra hyperparameters (e.g., the $k$ and $\lambda$ of RAPS).
@@ Appendix F Comparison with related libraries
Numerous libraries implement conformal prediction (CP) algorithms, yet their suitability for deep learning scenarios differs significantly, particularly in supporting modern deep learning models like GNNs and LLMs, as well as their scalability for large-scale datasets. In Table 8, we present a concise comparison of open-source CP libraries, focusing on supported model types and functionalities:
@@
In the following, we compare several popular CP libraries, emphasizing TorchCP’s superior compatibility and performance with deep learning applications.
nonconformist (Linusson et al., 2020): A NumPy-based library supporting basic CP algorithms for classification and regression tasks. However, it lacks GPU acceleration, batch processing, and compatibility with deep learning models like DNN Classifiers, DNN Regressors, GNNs, or LLMs, making it unsuitable for complex deep learning scenarios. While nonconformist implements predictors and scores as different classes following a low-coupling design, its algorithmic coverage is limited, implementing only inductive CP and aggregated CP with 5 scores.
@@
Crepes (Boström, 2022): a NumPy-based library with extensive prediction forms that supports the classifiers, regressors, and predictive systems with semi-online CP, excelling in statistical classification and regression tasks. However, it lacks GPU acceleration, batch processing, and seamless extensions to modern models like GNNs or LLMs, limiting scalability for engineers tackling large-scale or complex deep learning problems. Similar to TorchUQ, Crepes implements non-conformity scores as functions passed to predictors, so it also suffers from supporting scores with additional hyperparameters (e.g., $k$ and $\lambda$ in RAPS).
PUNCC (Mendil et al., 2023): A scikit-learn-compatible library delivers Conformal Prediction for tasks like regression, classification, and anomaly detection, but struggles with modern deep learning applications. It does not support advanced deep learning models such as DNNs, GNNs, or LLMs, lacks GPU acceleration and batch processing, and performs significantly slower on large datasets (up to 10x slower than TorchCP in ImageNet, see Figure 6(b)). PUNCC defines non-conformity scores as functions, but ties each score to a dedicated predictor, resulting in high coupling.